OpenStudy (anonymous):

how many moles are in 3.4*10^-7 grams of silicon dioxide

OpenStudy (anonymous):

all this is asking for is for you to convert your Amount in Grams to molar mass. So what you have to do is find the Compound first. Silicon is SI and dioxide is O2 so its SiO2 now what u do is find the molar mass of these elements. Si =28.08 O=16*2 because u have two atoms per molecule. So the total is about 60.02g of SiO2 for every mole. so now what u do is divide 3.4*10^7g*1/60.02g

OpenStudy (anonymous):

5.66*10^-9 of moles is your answer.

OpenStudy (rina.r):

Moles = mass / molar mass Molar mass SiO2 = 28 + 32 = 60 so moles sand = 3.4 x 10-7 / 60 http://answers.yahoo.com/question/index?qid=20101117113445AAnBF7o